Under Armor has a monopoly position in a certain type of track shoe. They face a specific cost function,

TC= $80,000+25Q

What is Under Armor’s marginal cost?

If the price elasticity of demand for the shoes is -1.75, what price should they charge?

What is the marginal revenue in part B?

If Nike enters the market for the same shoe and price elasticity increases to -2.5, what price should Under Armor charge?
Answer #1

TC= $80,000+25Q

MC= dTC/dQ = 25

we kow this condition

MR = P*(1-1/e), also we know that at profit maximising position the MR=MC

so, MR = MC = P*(1-1/e)

25 = P*(1-1/1.75)

P = 58.33 The MR will also be equal to MC = 25

Again for e = 2.5

MR = MC = P*(1-1/e)

25 = P*(1-1/2.5)

P = 41.67